# Automated Trustworthiness Oracle Generation for Machine Learning Text Classifiers

### v1

@@ Abstract.
Machine learning (ML) for text classification has been widely used in various domains, such as toxicity detection, chatbot consulting, and review analysis. These applications can significantly impact ethics, economics, and human behavior, raising serious concerns about trusting ML decisions. Several studies indicate that traditional uncertainty metrics, such as model confidence, and performance metrics, like accuracy, are insufficient to build human trust in ML models. These models often learn spurious correlations during training and predict based on them during inference. When deployed in the real world, where such correlations are absent, their performance can deteriorate significantly. To avoid this, a common practice is to test whether predictions are made reasonably based on valid patterns in the data, Along with this, a challenge known as the trustworthiness oracle problem has been introduced. So far, due to the lack of automated trustworthiness oracles, the assessment requires manual validation, based on the decision process disclosed by explanation methods. However, this approach is time-consuming, error-prone, and not scalable.
To address this problem, we propose TOKI, the first automated trustworthiness oracle generation method for text classifiers. TOKI automatically checks whether the words contributing the most to a prediction are semantically related to the predicted class. Specifically, we leverage ML explanation methods to extract the decision-contributing words and measure their semantic relatedness with the class based on word embeddings. As a demonstration of its practical usefulness, we also introduce a novel adversarial attack method that targets trustworthiness vulnerabilities identified by TOKI. We compare TOKI with a naive baseline based solely on model confidence. To evaluate their alignment with human judgement, experiments are conducted on human-created ground truths of approximately 6,000 predictions. Additionally, we compare the effectiveness of TOKI-guided adversarial attack method with A2T, a state-of-the-art adversarial attack method for text classification. Results show that (1) relying on prediction uncertainty metrics, such as model confidence, cannot effectively distinguish between trustworthy and untrustworthy predictions, (2) TOKI achieves 142% higher accuracy than the naive baseline, and (3) TOKI-guided adversarial attack method is more effective with fewer perturbations than A2T.
Warning: This paper contains extracts from datasets on mental health and hate speech, which may upset some readers. Reader discretion is advised.
@@ 1. Introduction
Machine learning (ML) holds significant importance in contemporary advanced systems, such as spam detection, clinical text analysis, and vulnerability detection, with text classification being a primary application. Despite their superior performance during development, ML models can still fail in real-life scenarios (Caruana et al., 2015; Lapuschkin et al., 2019), raising concerns about trusting their decisions. When assessing a model, it is important not only to evaluate its general task-solving ability using metrics, such as prediction uncertainty or classification accuracy, but also to understand its decision-making process (Liao and Wortman Vaughan, 2024). Indeed, several studies show that these metrics alone are insufficient indicators of model reliability (Nguyen et al., 2015; Canbek et al., 2022),
@@
An ML model relying on irrelevant features, when faced with unseen data, can misclassify in the absence of these features or be fooled by them, leading to problematic predictions. We illustrate this using a Bert-based binary sentiment classifier (Devlin et al., 2019) applied to amazon reviews. Figure 1(d) shows four review examples, the model’s predictions, and the probabilities of each class. We also highlight the important words steering these predictions, which are identified by LIME (Ribeiro et al., 2016), an explanation method. Orange and blue colors highlight words contributing to the negative and positive classes, respectively, with darker shades indicating higher importance. The review in Figure 1(a) is correctly classified as negative with high confidence. However, a closer look at the most important words the classifier relies on reveals that words playing a vital role, such as “back”, “the”, “anything”, and “through”, are unrelated to either class. When removing some of them, as shown in Figure 1(b), the review is misclassified as positive by the same model, despite negative sentiment in phrases “not seen him CLEAN”, “slow”, “rough patch”, and “sloggin”. To further explore the potential harm of this phenomenon, we synthesise a new review using a positive review shown in Figure 1(c). This new review, as illustrated in Figure 1(d), is injected with the words “back”, “through”, and “anything”. Although, these injections do not change the original meaning, the confident correct prediction is flipped to its opposite. These shifts underscore the importance of assessing whether the learned patterns are genuinely valid and generalisable or if they are merely based on spurious correlations within the training data, typically referred to as shortcut learning (Geirhos et al., 2020; Du et al., 2023). A shortcut learning model is unlikely to provide correct classifications for the right reasons. Therefore, it becomes ineffective once deployed in the real world, where spurious correlations are absent.
@@
To ensure the quality of ML systems in a reliable and cost-effective way, considerable effort has been focused on automating various aspects of the testing process, particularly through automated test oracles. Traditional software testing has inspired automated test oracles for ML to test various properties, such as correctness, fairness, and robustness using techniques like metamorphic testing (Pei et al., 2017; Guo et al., 2018). However, trustworthiness testing has not received as much attention and consequently remains less resolved. Trustworthiness refers to the ability of a model to make reasonable predictions based on relevant features, such as semantically related words in text classification. Developing automated test oracles (Barr et al., 2015) for trustworthiness testing is a long-lasting challenge hindering the broader adoption of ML models in the real world. Due to the lack of automated trustworthiness oracles, most studies (Ribeiro et al., 2016; Lapuschkin et al., 2019; Du et al., 2021) rely on human-based evaluation as the oracle to assess the ML prediction reasoning uncovered by explanation methods (Zhao et al., 2024). However, these approaches are time-consuming, error-prone, and not scalable (Lertvittayakumjorn and Toni, 2021; Ye et al., 2024).
@@
A benchmark for trustworthiness assessment for text classification, containing approximately 6,000 predictions in various domains, including topic classification, sentiment analysis, clinical mental text classification, and hate speech detection.
@@
An ablation study and comparative evaluations of TOKI. For trustworthiness assessment, we compare TOKI with a naive baseline solely based on model confidence. For adversarial attacks, we compare our method with A2T (Yoo and Qi, 2021), a state-of-the-art adversarial attack method. The results demonstrate TOKI’s superior effectiveness and efficiency.
2. Problem Definition and Motivation
@@ 2.1. Definition of Trustworthiness
Trustworthiness is a complex concept that has raised numerous scholarly debates among researchers. This paper focuses on trustworthiness, in particular, whether a model makes predictions based on valid and reasonable patterns rather than spurious ones. Hence, we adopt the definition of trustworthiness proposed by Kästner et al. (Kästner et al., 2021) as shown in Definition 1.
@@
Kästner et al. also distinguish between trustworthiness and trust: trustworthiness is a property of an ML model, while trust is the perception a person has towards it (Kästner et al., 2021). Therefore, people can still trust an untrustworthy ML model.
According to Definition 1, understanding a model thoroughly helps justify our beliefs about how well it works. Model explanations can serve as a means to gain this understanding and justification (Wiegreffe and Pinter, 2019), thereby promoting trustworthiness (Bussone et al., 2015; Kästner et al., 2021). There are two main types of model explanations: global (Caruana et al., 2015), which provides insights into the entire model’s inner workings, and local (Ribeiro et al., 2016; Li et al., 2016; Mohebbi et al., 2021), which focuses on individual predictions. Local explanations, by breaking down a model into its components, allow users to grasp its functionality and decision-making process in a way that aligns with human cognitive patterns. For this reason, they are more readily applicable (Adadi and Berrada, 2018). Hence, we leverage local explanations to uncover the reasoning behind individual predictions by highlighting the most relevant features contributing to those predictions.
A trustworthy model should make correct predictions, and the reasoning behind them, supported by explanations, should also be plausible. In this way, we propose the following definition of a trustworthy prediction.
@@
In the context of text classification, a trustworthy prediction should rely on words in the input text that align with human reasoning. To simulate human judgement, the reasoning behind a prediction can be considered plausible if words contributing the most to the prediction are semantically related to the predicted class. We follow the definition of semantic relatedness described by Budanitsky et al. (Budanitsky and Hirst, 2006). Words can be semantically related by lexical relationships, such as meronymy (car–wheel) and antonymy (hot–cold), or just by any kind of functional association or other “non-classical relations” (pencil–paper, penguin–Antarctica, and rain–flood) (Morris and Hirst, 2004). It is important to distinguish semantic relatedness from semantic similarity. Semantic relatedness is a more general concept (Budanitsky and Hirst, 2006) while semantic similarity refers to the degree of overlap or resemblance in meaning between two words (Slimani, 2013). For example, “desk” and “chair” are semantically related but not semantically similar, “desk” and “table”, on the other hand, are both semantically related and semantically similar. It is also important to note that semantically relatedness is contextual, meaning two words are semantically related in a specific context, but might not in others. For instance, “bank” in “the bank exploits small firms” is semantically related to economics, while “bank” in “we walked along the river bank” is not.
@@ 2.2. Trustworthiness and Robustness
Sharing the same goal of improving the model’s generalisability, robustness testing might be able to identify problems that can be spotted during trustworthiness testing. Local robustness measures the model’s ability to retain its prediction on a sample under perturbations, also known as adversarial examples (Zhang et al., 2020b), that do not affect human perception and decision. For example, removing the header in Figure 1(b) should not alter a human’s decision to classify it as Atheism. However, the model fails to maintain its original decision on this example, and we can conclude that the adversarial example reveals a local robustness issue.
@@
We differentiate trustworthiness from robustness by revisiting the prediction in Figure 1(a), showing that robustness testing is unable to identify issues in the model’s behavior, but trustworthiness testing can. TextAttack, a framework implementing several state-of-the-art adversarial example generation methods (Morris et al., 2020), is applied to attack (Zhang et al., 2020b) the model on the same text input. An additional adversarial example is generated manually to ensure grammar and semantic preservation by replacing orange-highlighted words in Figure 1(a) with their synonyms from WordNet (Miller, 1995). The substitutions include {① idea–recommendation, ② BUT–NEVERTHELESS, ③ It–This, ④ the–this, ⑤ back–end, ⑥ anything–whatsoever, ⑦ slow–obtuse, ⑧ I–me, ⑨ through–by, ⑩ not–n’t}. Figures 2(a) and 2(b), with bold words representing perturbations, show LIME explanations for two predictions to adversarial examples generated by TextAttack and manually, respectively. All adversarial examples fail to attack the model since it maintains the same prediction. These results indicate that although the model is robust against certain adversarial examples, its prediction can still be untrustworthy.
@@ 2.3. Trustworthiness Oracle
ML testing involves providing a model with inputs and observing its responses. The oracle problem (Barr et al., 2015) in ML testing is the challenge of determining whether these responses are appropriate. In trustworthiness testing, a response is an explainable prediction, as described in Definition 3.
@@
Three cases can occur with an explainable prediction: (1) incorrect, (2) correct due to semantically related words, and (3) correct due to semantically unrelated words. In the first case, investigating incorrectness becomes more important than trustworthiness assessment. Although the explanation might hint at why the prediction is incorrect, we leave this interesting avenue for future work. In this way, this paper only focuses on differentiating the last two cases, meaning that it examines only correct explainable predictions in the context of trustworthiness testing. We then adopt the definition of test oracles from Barr et al. (Barr et al., 2015). Definition 4 describes trustworthiness test data while Definition 5 defines trustworthiness oracles.
@@ 2.4. Applications of Trustworthiness Oracles
The trustworthiness problem exists in various ML applications. For example, a clinical mental text classifier from social media posts shows good held-out performance but might rely on irrelevant clues frequently present in the training data, such as post tags, job positions, and special occasions (Harrigian et al., 2020). When deployed in the real world, such classifier is unlikely to perform well on unseen data. Hence, it is crucial to assess the reasoning of ML models to detect any behavioral issues before deploying them in real-world scenarios, rather than only evaluating their classification performance using metrics such as accuracy, precision, recall, and F1-score.
@@
Figure 3(b) shows the process integrated with trustworthiness testing. In addition to evaluating traditional performance metrics, this process also assesses the model’s behavior behind correct predictions. In this process, trustworthiness oracles determine whether a prediction is trustworthy or based on spurious correlations. Identifying such issues allows the engineers to mitigate the impact of spurious correlations, ensure the model is correct for the right reasons, and improve its generalizability. If undetected, these issues can lead to problematic predictions in the future. Human annotations often serve as trustworthiness oracles, but this approach is not scalable (Ye et al., 2024). For this reason, developing automated trustworthiness oracles can greatly advance the development and applications of ML systems in the real world.

### v3

@@ Abstract.
Machine learning (ML) for text classification has been widely used in various domains, such as toxicity detection, chatbot consulting, and review analysis. These applications can significantly impact ethics, economics, and human behavior, raising serious concerns about trusting ML decisions. Several studies indicate that traditional uncertainty metrics, such as model confidence, and performance metrics, like accuracy, are insufficient to build human trust in ML models. These models often learn spurious correlations during training and predict based on them during inference. When deployed in the real world, where such correlations are absent, their performance can deteriorate significantly. To avoid this, a common practice is to test whether predictions are made reasonably based on valid patterns in the data. Along with this, a challenge known as the trustworthiness oracle problem has been introduced. So far, due to the lack of automated trustworthiness oracles, the assessment requires manual validation, based on the decision process disclosed by explanation methods. However, this approach is time-consuming, error-prone, and not scalable.
To address this problem, we propose TOKI, the first automated trustworthiness oracle generation method for text classifiers. TOKI automatically checks whether the words contributing the most to a prediction are semantically related to the predicted class. Specifically, we leverage ML explanation methods to extract the decision-contributing words and measure their semantic relatedness with the class based on word embeddings. As a demonstration of its practical usefulness, we also introduce a novel adversarial attack method that targets trustworthiness vulnerabilities identified by TOKI. We compare TOKI with a naive baseline based solely on model confidence. To evaluate their alignment with human judgement, experiments are conducted on human-created ground truths of approximately 8,000 predictions. Additionally, we compare the effectiveness of TOKI-guided adversarial attack method with A2T, a state-of-the-art adversarial attack method for text classification. Results show that (1) relying on prediction uncertainty metrics, such as model confidence, cannot effectively distinguish between trustworthy and untrustworthy predictions, (2) TOKI achieves 142% higher accuracy than the naive baseline, and (3) TOKI-guided adversarial attack method is more effective with fewer perturbations than A2T.
Warning: This paper contains examples of language that some people may find offensive or upsetting.
@@ 1. Introduction
Machine learning (ML) holds significant importance in contemporary advanced systems, such as spam detection, clinical text analysis, and vulnerability detection, with text classification being a primary application. Despite their superior performance during development, ML models can still fail in real-life scenarios (Caruana et al., 2015; Lapuschkin et al., 2019), raising concerns about trusting their decisions. When assessing a model, it is important not only to evaluate its general task-solving ability using metrics, such as prediction uncertainty or classification accuracy, but also to understand its decision-making process (Liao and Wortman, 2024). Indeed, several studies show that these metrics alone are insufficient indicators of model reliability (Nguyen et al., 2015; Canbek et al., 2022),
@@
An ML model relying on irrelevant features, when faced with unseen data, can misclassify in the absence of these features or be fooled by them, leading to problematic predictions. We illustrate this using a Bert-based (Devlin et al., 2019) binary sentiment classifier applied to amazon reviews. Figure 1(d) shows four review examples, the model’s predictions, and the probabilities of each class. We also highlight the important words steering these predictions, which are identified by LIME (Ribeiro et al., 2016), an explanation method. Orange and blue colors highlight words contributing to the negative and positive classes, respectively, with darker shades indicating higher importance. The review in Figure 1(a) is correctly classified as negative with high confidence. However, a closer look at the most important words the classifier relies on reveals that these words, such as “back”, “the”, “anything”, and “through”, are unrelated to either class. When removing some of them, as shown in Figure 1(b), the review is misclassified as positive by the same model, despite negative sentiment in phrases “too slow”, “rough patch”, “sloggin”, and “give up”. To further explore the potential harm of this phenomenon, we synthesise a new review using a positive review shown in Figure 1(c). This new review, as illustrated in Figure 1(d), is injected with the words “back”, “through”, and “anything”. Although, these injections do not change the original meaning, the confident correct prediction is flipped to its opposite. These shifts underscore the importance of assessing whether the learned patterns are genuinely valid and generalisable or if they are merely based on spurious correlations within the training data, typically referred to as shortcut learning (Du et al., 2023). A shortcut learning model is unlikely to provide correct classifications for the right reasons. Therefore, it becomes ineffective once deployed in the real world, where spurious correlations are absent.
@@
To ensure the quality of ML systems in a reliable and cost-effective way, considerable effort has been focused on automating various aspects of the testing process, particularly through automated test oracles (Barr et al., 2015). Traditional software testing has inspired automated test oracles for ML to test various properties, such as correctness, fairness, and robustness using techniques like metamorphic testing (Pei et al., 2017; Guo et al., 2018). However, trustworthiness testing has not received as much attention and consequently remains less resolved (Cho et al., 2024). Trustworthiness refers to the ability of a model to make reasonable predictions based on relevant features, such as semantically related words in text classification. Developing automated test oracles for trustworthiness testing is a long-lasting challenge hindering the broader adoption of ML models in the real world. Due to the lack of automated trustworthiness oracles, most studies (Ribeiro et al., 2016; Lapuschkin et al., 2019; Du et al., 2021) rely on human-based evaluation as the oracle to assess the ML prediction reasoning uncovered by explanation methods (Zhao et al., 2024). However, this is time-consuming, error-prone, and not scalable (Ye et al., 2024).
@@
A benchmark for trustworthiness assessment for text classification, which contains approximately 8,000 predictions in various domains, such as topic classification, sentiment analysis, clinical mental text classification, hate speech detection, and software issue management.
@@
Ablation studies and comparative evaluations of TOKI. For trustworthiness assessment, we compare TOKI with a naive baseline solely based on model confidence. For adversarial attacks, we compare our method with A2T (Yoo and Qi, 2021), a state-of-the-art (SOTA) adversarial attack method. The results show TOKI’s superior effectiveness and efficiency.
2. Problem Definition, Preliminary, and Motivation
@@ 2.1. Definition of Trustworthiness
Trustworthiness is a complex concept that has raised numerous scholarly debates among researchers. This paper focuses on trustworthiness, in particular, whether a model makes predictions based on valid and reasonable patterns rather than spurious ones. Hence, we adopt the definition of trustworthiness proposed by Kästner et al. (2021) as shown in Definition 1.
@@
They also distinguish trustworthiness from trust: trustworthiness is a system’s property, as recognised by prior studies as a critical non-functional requirement (Riccio et al., 2020), whereas trust is the perception a person has towards it. Hence, people can still trust an untrustworthy system.
According to Definition 1, understanding a model thoroughly helps justify our beliefs about how well it works. Model explanations can serve as a means to gain this understanding (Wiegreffe and Pinter, 2019), thereby promoting trustworthiness (Bussone et al., 2015; Kästner et al., 2021). There are two main types of model explanations: global (Caruana et al., 2015), which provides insights into the entire model’s inner workings, and local (Ribeiro et al., 2016; Li et al., 2016; Mohebbi et al., 2021), which focuses on individual predictions. Local explanations, by breaking down a model into its components, allow users to grasp its functionality and decision-making process in a way that aligns with human cognitive patterns. Hence, they are more readily applicable (Adadi and Berrada, 2018). In text classification, various local explanations have been developed, such as feature attribution-based (Li et al., 2016; Mohebbi et al., 2021), attention-based (Yeh et al., 2024; Barkan et al., 2021), and counterfactual (Treviso et al., 2023) explanations. While other local explanations are subject to extensive debate (Jain and Wallace, 2019), feature attribution-based explanations, such as LIME (Ribeiro et al., 2016), have proven to be effective, faithful, and widely used (Mariotti et al., 2024). Hence, we leverage these explanations to uncover the reasoning behind individual predictions by highlighting the most relevant features contributing to those predictions. Exploring alternative explanations for assessing trustworthiness is an interesting avenue for future work.
A trustworthy model should make correct predictions, and the reasoning behind them, supported by explanations, should also be plausible. We propose the definition of a trustworthy prediction.
@@
In the context of text classification, a trustworthy prediction should rely on words in the input text that align with human reasoning. To simulate human judgement, the reasoning behind a prediction can be considered plausible if words contributing the most to the prediction are semantically related to the predicted class. We follow the definition of semantic relatedness described by Budanitsky and Hirst (2006). Words can be semantically related by lexical relationships, such as meronymy (car–wheel) and antonymy (hot–cold), or just by any kind of functional association or other “non-classical relations” (pencil–paper, penguin–Antarctica, and rain–flood) (Morris and Hirst, 2004). It is important to distinguish semantic relatedness from semantic similarity. Semantic relatedness is a more general concept (Budanitsky and Hirst, 2006) while semantic similarity refers to the degree of overlap or resemblance in meaning between two words (Slimani, 2013). For example, “desk” and “chair” are semantically related but not semantically similar, “desk” and “table”, on the other hand, are both semantically related and semantically similar. It is also important to note that semantically relatedness is contextual, meaning two words are semantically related in a specific context, but might not in others. For instance, “bank” in “the bank exploits small firms” is semantically related to economics, while “bank” in “we walked along the river bank” is not.
@@ 2.2. Trustworthiness and Robustness
Sharing the same goal of improving the model’s generalisability, robustness testing might be able to identify problems that can be spotted during trustworthiness testing. Local robustness measures the model’s ability to retain its prediction on a sample under perturbations, also known as adversarial examples (Zhang et al., 2020b), that do not affect human perception and decision. For example, replacing a few words in Figure 1(a) with their synonyms should not alter a human’s decision to classify it as negative. If the model fails to maintain its original prediction on this adversarial example, we can conclude that the adversarial example reveals a local robustness issue.
@@
We differentiate trustworthiness from robustness by revisiting the prediction in Figure 1(a), showing that robustness testing is unable to identify issues in the model’s behavior, but trustworthiness testing can. TextAttack (Morris et al., 2020), a framework implementing several SOTA adversarial example generation methods, is applied to attack (Zhang et al., 2020b) the model on the same text input. An additional adversarial example is generated manually to preserve grammar and semantic by replacing orange-highlighted words in Figure 1(a) with their synonyms from WordNet (Miller, 1995). The substitutions are ① idea–recommendation, ② BUT–NEVERTHELESS, ③ It–This, ④ the–this, ⑤ back–end, ⑥ anything–whatsoever, ⑦ slow–obtuse, ⑧ I–me, ⑨ through–by, ⑩ not–n’t. Figures 2(a) and 2(b), with bold words representing perturbations, show LIME explanations for two predictions to adversarial examples generated by TextAttack and manually, respectively. All adversarial examples fail to attack the model since it maintains the same prediction. This indicates that although the model is robust against certain adversarial examples, its prediction can still be untrustworthy.
@@ 2.3. Trustworthiness Oracle
ML testing involves providing a model with inputs and observing its responses. The oracle problem in ML testing is the challenge of determining whether these responses are appropriate. In trustworthiness testing, a response is an explainable prediction, as described in Definition 3.
@@
Three cases can occur with an explainable prediction: (1) incorrect, (2) correct due to semantically related words, and (3) correct due to semantically unrelated words. In the first case, investigating incorrectness becomes more important than trustworthiness assessment. Although the explanation might hint at why the prediction is incorrect, we leave this interesting avenue for future work. In this way, this paper only focuses on differentiating the last two cases, meaning that it examines only correct explainable predictions in the context of trustworthiness testing. We then adopt the definition of test oracles from Barr et al. (2015). Definition 4 describes trustworthiness test data while Definition 5 defines trustworthiness oracles.
@@ 2.4. Applications of Trustworthiness Oracles
The trustworthiness problem exists in various ML applications. For example, a clinical mental text classifier from social media posts shows good held-out performance but might rely on irrelevant clues frequently present in the training data, such as post tags, job positions, and special occasions (Harrigian et al., 2020). When deployed in the real world, such classifier is unlikely to perform well on unseen data. Hence, it is crucial to assess the reasoning of ML models to detect any behavioral issues before deploying them in real-world scenarios, rather than only evaluating their classification performance using metrics such as accuracy, recall, and F1-score.
Figure 3(b) shows the process integrated with trustworthiness testing. In addition to evaluating traditional performance metrics, this process also assesses the model’s behavior behind correct predictions. In this process, trustworthiness oracles determine whether a prediction is trustworthy or based on spurious correlations. Identifying such issues allows the engineers to mitigate the impact of spurious correlations, ensure the model is correct for the right reasons, and improve its generalizability. If undetected, these issues can lead to problematic predictions in the future. Human annotations often serve as trustworthiness oracles, but this approach is not scalable (Ye et al., 2024). Integrating automated trustworthiness oracles into the software engineering (SE) lifecycle for ML systems can greatly advance their development and applications. Automated trustworthiness oracles enable the automation of trustworthiness testing, which is vital and closely intertwined with other SE activities (Riccio et al., 2020). This is especially valuable in the iterative development of ML systems where performance and trustworthiness must be evaluated and refined (Martínez et al., 2022). Specifically, automated trustworthiness oracles support continuous integration and delivery pipelines by automating trustworthiness testing for ML systems. They also serve as a tool for monitoring ML systems in real-world environments, particularly in online testing or DevOps workflows, to verify whether predictions are trustworthy in real time, and test with real-world and corner-cased inputs. Moreover, the outputs of trustworthiness oracles provide actionable insights for feedback-driven repairing and improving processes, reducing the need for human intervention.
@@
In addition, trustworthiness oracles help assess the trustworthiness of systems in SE that employ ML for text classification. ML has been widely used in SE for text classification, automating numerous tasks to enhance planning, design, and maintenance. For example, sentiment analysis has been applied to various SE artifacts, including git commit comments (Sinha et al., 2016), JIRA issues (Ortu et al., 2015), and apps’ reviews (Panichella et al., 2015). It also helps assess developers’ psychological states (Guzman and Bruegge, 2013), and analyse sentiment on Q&A sites like StackOverflow to recommend improvements for source code (Rahman et al., 2015) or to identify problematic API design features (Zhang and Hou, 2013). In addition to sentiment analysis, other tasks, such as software requirement classification (Pérez-Verdejo et al., 2020) and project issue categorisation (Schulte et al., 2024), have also been integrated with ML to reduce human effort and improve efficiency. Hence, automated trustworthiness oracles offer opportunities to foster greater trust in these systems, both within SE domains and more broadly.